OpenStudy (shal):

help with algebra 2? will give medal and become fan please help

OpenStudy (smurfy14):

type your question

OpenStudy (shal):

check yur messages

OpenStudy (happykiddo):

Dude... Someone would have already helped you if you just posted your question, instead of sending out, Can someone help me?

OpenStudy (shal):

@newtonson

OpenStudy (shal):

@izzy100200

OpenStudy (shal):

@jabez177

OpenStudy (shal):

help please

OpenStudy (anonymous):

THE ANSWER OF YOUR QUESTION IS -(5/2)

jabez177 (jabez177):

Okay. We need to find the slope so do you remember which formula we are going to use?

OpenStudy (smurfy14):

When given two points on a graph to find the slope between them first count the distance between the two horizontally (left or down directions are negative) and then go vertically. So in this case answer is -5/2

OpenStudy (anonymous):

 CONSIDER THESE TWO POINTS AND FIND THE SLOPE USING FORMULA Y2-Y1/X2-X1.

jabez177 (jabez177):

We are going to use the formula: \[\frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (smurfy14):

Cause you go left 5 (negative 5) and then up two (up and right directions are positive).

OpenStudy (shal):

thank you

jabez177 (jabez177):

So I don't have to explain anymore? :P

OpenStudy (shal):

slope help please?

OpenStudy (shal):

@newtonson

OpenStudy (shal):

what is the slope of 2x+y=4

OpenStudy (shal):

slope of the line*

OpenStudy (shal):

@Lovelarap

OpenStudy (anonymous):

if i'm not wrong, the slope should be -2 because you take 2x to the other side and the number in front of the x is the slope

OpenStudy (shal):

a line has a slope of -6 that passes through the point (-12,5) what is the equation of the line?

OpenStudy (shal):

help please

OpenStudy (anonymous):

y=-6x+b 5=-6(12)+b b=77 y=-6x+77 I think

OpenStudy (shal):

choices are 6x+y=-77 -6y+y=77 -6+y=17 6x+y=67

OpenStudy (shal):

-67*

OpenStudy (anonymous):

6x+y=77

OpenStudy (shal):

thts not a choice

OpenStudy (anonymous):

D

OpenStudy (shal):

@mustbethestars

OpenStudy (shal):

last 3